# Key Words.: binaries: spectroscopic – binaries: close – stars: early-type – stars: massive – Magellanic Clouds1 Introduction

### v1

@@
Binarity at LOw Metallicity (BLOeM):††thanks: Based on observations collected at the European Southern Observatory under ESO program ID 112.25W2.
Early B-type stars with initial masses between 8 and 15 are frequently found in multiple systems, as evidenced by multi-epoch spectroscopic campaigns in the Milky Way and the Large Magellanic Cloud (LMC). Previous studies have shown no strong metallicity dependence in the close-binary ( au) fraction or orbital-period distributions between the Milky Way’s solar metallicity () and that of the LMC (). However, similar analyses for a large sample of massive stars in more metal-poor environments are still scarce. We focus on 309 early B-type stars (luminosity classes III–V) from the Binarity at LOw Metallicity (BLOeM) campaign, which targeted nearly 1000 massive stars in the Small Magellanic Cloud (SMC, ) using VLT/FLAMES multi-epoch spectroscopy. By applying binary detection criteria consistent with previous works on Galactic and LMC samples, we identify 153 stars (91 SB1, 59 SB2, 3 SB3) exhibiting significant radial-velocity (RV) variations, resulting in an observed multiplicity fraction of % . Using Monte Carlo simulations to account for observational biases, we infer an intrinsic close-binary fraction of % . This fraction reduces to % when increasing our RV threshold from 20 to 80 km s-1, however, an independent Markov chain Monte Carlo analysis of the peak-to-peak RV distribution () confirms a high multiplicity fraction of % . These findings suggest a possible anti-correlation between metallicity and the fraction of close B-type binaries, with the SMC multiplicity fraction significantly exceeding previous measurements in the LMC and the Galaxy. The enhanced fraction of close binaries at SMC’s low metallicity may have broad implications for massive-star evolution in the early Universe. More frequent mass transfer and envelope stripping could boost the production of exotic transients, stripped supernovae, gravitational-wave progenitors, and sustained UV ionising flux, potentially affecting cosmic reionisation. Theoretical predictions of binary evolution under metal-poor conditions will provide a key test of our results.
@@
Massive stars go through complex evolutionary scenarios that are significantly influenced by their high multiplicity fraction (Mason et al., 2009; Chini et al., 2012; Sana et al., 2014; Kobulnicky et al., 2014; Dunstall et al., 2015; Moe & Di Stefano, 2017; Ramírez-Tannus et al., 2024) and frequent binary interactions (Sana et al., 2012; de Mink et al., 2014). Spectroscopic studies in the Galaxy and the Large Magellanic Cloud (LMC) consistently indicate that more than 50% of O- and B-type stars are part of binary or multiple systems within the orbital period range where interactions are expected ( d). Including longer orbital periods, the overall multiplicity fraction can exceed 90% for the O-type stars (Sota et al., 2014; Bordier et al., 2022). Sana et al. (2012) found an intrinsic multiplicity fraction (after correction for observational biases) of % for a sample of 71 O-type stars in young Galactic clusters. In the Cygnus OB2 association, Kobulnicky et al. (2014, and references therein) presented orbital solutions for 48 massive stars with spectral types ranging from O5 to B2, resulting in an intrinsic multiplicity fraction near 55% from a sample of 45 O-type and 83 B-type stars. More recently, Banyard et al. (2022) studied the B-type content of the young Galactic cluster NGC 6231, determining % from 80 stars covering the entire B-type spectral type range (B0-B9). While further multiplicity studies have been conducted in the Milky Way (MW), many have not been corrected for observational biases (e.g. Barbá et al., 2017; Berlanas et al., 2020; Ritchie et al., 2022), have targeted narrow spectral ranges (Abt et al., 1990), focused on peculiar types (e.g., chemically peculiar stars; Schöller et al., 2010), or addressed larger separations (Mason et al., 2009; Sana et al., 2014). These factors hinder direct comparison of multiplicity fractions due to the different observational strategies and biases.
@@
In the LMC (), the VLT-FLAMES Tarantula Survey (VFTS; Evans et al., 2011) performed multiplicity studies on 360 O-type (Sana et al., 2013) and 408 B-type stars (Dunstall et al., 2015, of which 361 were main-sequence and giant stars), resulting in binary fractions of % and % , respectively. The consistency across these studies suggests that the multiplicity properties of massive stars may be independent of metallicity. Furthermore, the distribution of orbital periods is remarkably similar across most of these samples (Almeida et al., 2017; Villaseñor et al., 2021; Banyard et al., 2022), suggesting that O- and B-type stars in multiple systems are formed with similar efficiency and orbital configurations within the metallicity ranges studied so far, underscoring the need for multiplicity studies of massive stars in lower metallicity environments.
@@
In the case of solar-type binaries, a strong anti-correlation between binary fraction and metallicity has been found (Badenes et al., 2018; Moe et al., 2019; Offner et al., 2023). Moe et al. (2019) computed intrinsic binary fractions for five different observed samples comprising spectroscopic and eclipsing binaries, and found a combined close binary fraction that drops from % at [Fe/H] to % at [Fe/H]. This stands in stark contrast to the findings for OB-type stars (see also Moe & Di Stefano, 2013) and also to solar-type stars in wider orbits ( au; El-Badry & Rix, 2019). Both Moe et al. (2019) and El-Badry & Rix (2019) agree that the formation of wide binaries is governed by turbulent core fragmentation which is independent of metallicity, but, in the case of close binaries, the main formation mechanism is disk fragmentation up to separations of au. Moe et al. (2019) suggest that the difference between the sharp increase in the binary fraction of solar-type stars towards low metallicities and the close-to-constant fraction observed in massive stars lies in the disk of massive protostars. Already at solar metallicity, these disks are already highly unstable and have a high probability of fragmentation, leaving little room to further increase multiplicity at reduced metallicities.
@@
The findings in the MW and LMC suggest a universality in the multiplicity properties of massive stars, and motivate further investigation on whether trends persist at lower metallicity. Since the LMC metallicity is only half solar, it is imperative to perform multiplicity studies of massive stars at even lower metallicity. In this context, the Small Magellanic Cloud (SMC) is a unique environment due to its proximity (62 kpc; Graczyk et al., 2020), low metallicity () representative of the high-redshift universe (Nakajima et al., 2023), and high content of massive stars (Evans et al., 2004). Furthermore, theoretical predictions indicate that the winds of massive stars at SMC metallicity are strongly reduced (Vink et al., 2001; Björklund et al., 2021), with empirical evidence showing that they might be even weaker than predicted (Ramachandran et al., 2019; Rickard et al., 2022). If there is a connection between metallicity and the multiplicity properties of massive stars, the SMC offers a prime environment for testing the universality of these properties. For other low-metallicity galaxies in the Local Group, resolving a significant sample of massive stars with the current generation of telescopes is not possible due to their greater distances.
@@
The Binarity at Low Metallicity (BLOeM; Shenar et al., 2024) campaign addresses this knowledge gap by focusing on the multiplicity properties of massive stars () within the SMC. BLOeM has observed 929 stars in eight different regions of the SMC. The sample has been divided into sub-samples based on their spectral types, emission properties and luminosity classes, enabling an interpretation of the results within the context of each sub-sample. The five samples are: the O-type stars (Sana et al., 2024); the B-type supergiants (spectral types B0-B3, Britavskiy et al., 2025); the B-type dwarfs and giants (this work); the cooler supergiants (spectral types later than B3, Patrick et al., 2025); and the emission-line stars (Oe/Be, Bodensteiner et al., 2025).
@@
In this study, we present the first detailed analysis of the multiplicity properties of a large sample of early B-type dwarf and giant stars at the low metallicity of the SMC. In Sect. 2, we introduce the sample of B-type stars and their main properties. Section 3 details our methodology to measure radial velocities and our criteria to classify spectroscopic binaries. Our results on the multiplicity properties are presented in Sec. 4, and our search for periodicity is described in Sec. 5. We discuss our results in a broader context in Sect. 6, and finally provide our conclusions in Sect. 7.
@@ 2 The sample
The B-type dwarfs and giants sample comprises 309 stars with spectral types ranging from B0 to B2.5 and luminosity classes V to III111We note that O-type primaries with B-type companions are part of the O-type sample (Sana et al., 2024).. Spectral types for the full sample were presented in Shenar et al. (2024). Specifically, our sample consist of all early B-type “non-supergiant” objects, including main-sequence (MS, class V), subgiant (class IV), and relatively unevolved giant stars (class III). We expect our sample to primarily consists of MS stars undergoing core-hydrogen burning and stars that have just recently exhausted hydrogen in their cores. Bright giants and supergiants (classes II and I) form a separate early B-type supergiants subsample, which is the focus of Britavskiy et al. (2025). Hereafter, we refer to our sample as the III/V sample. This classification aligns with that used for the 30 Dor B-type stars, where Dunstall et al. (2015) defined “unevolved” stars as those with surface gravity dex and analysed their multiplicity properties separately from the supergiant sample. However, unlike the 30 Dor unevolved sample, our sample is primarily composed by giant stars (see Sect. 4.1). In the case of Be stars, we assume that most of them are binary interaction products (Bodensteiner et al., 2020b; Dallas et al., 2022; Dodd et al., 2024), therefore they will be the subject of a separate work (Bodensteiner et al., 2025).
BLOeM has currently obtained and reduced nine epochs from the first semester of observations using the FLAMES instrument (Fibre Large Array Multi Element Spectrograph; Pasquini et al., 2002) in GIRAFFE mode () on the UT2 unit telescope of the Very Large Telescope (VLT) at the European Southern Observatory (ESO) in Chile. For ten stars in our sample the number of epochs is below nine, fluctuating between three and seven epochs, and in five additional stars one of the epochs has a too low signal-to-noise ratio (), thus only eight epochs were used. Additional epochs are expected by the end of 2024, with a total of 25 epochs planned by the conclusion of the campaign in September 2025. For a complete overview of the sample selection, data reduction, and spectral classification, see Shenar et al. (2024).
Each field’s central coordinates and the modified Julian dates (MJDs) per observation are provided by Shenar et al. (2024, Table A.1). Here, we present the total baseline of observations and minimum cadence for each field in Table 1. We have a mean observational baseline of 45 d and a minimum cadence of 1 to 2 d. This cadence allows us to constrain most short-period orbits, except for unevolved contact binaries with periods of less than d. However, these are likely to be eclipsing binaries and therefore detectable by the Optical Gravitational Lensing Experiment (OGLE; Pawlak et al., 2016) through their light curves (see Sect. 5.1).
@@
The distribution of spectral types per field is shown in Fig. 2. The overall distribution is very similar to that found by Ramachandran et al. (2019) in the supergiant shell SMC-SGS1 in the Wing of the SMC, a region that was not covered by BLOeM. Our sample is dominated by spectral types B1.5 and B2, corresponding roughly to masses between 9 and 12. There is a good range of spectral types across different fields, but in fields 7 and 8, the number of B1.5 and B2 seem to be overabundant with respect to the earliest stars (B0–B0.2), potentially suggesting an older population. We computed the ratio of latest-to-earliest type stars for these spectral types (i.e. (B1.5–B2)/(B0–B0.2)), finding ratios of 4.4 and 18.0 for fields 7 and 8 respectively, and a mean of 2.9 for the other six fields. Performing a Z-test returns that these differences are significant, with a z-score of 5 (field 7) and 49 (field 8) and p-values of (field 7) and 0.0 (field 8). This is consistent with fields 7 and 8 being areas without prominent star-forming regions (see Fig. 1 and details in main text). In contrast, field 3, which exhibits several H ii regions, has the lowest ratio of latest-to-earliest stars with 1.5. We note, however, that an uncertainty of one spectral subtype should be expected. Such uncertainty could at least partially explain the lack of B0.2 and B0.5 stars. These are stars that show weak He ii, and since He ii 4686 (stronger than He ii 4542) is not covered by the FLAMES LR02 setup, they are more difficult to classify. Given the small number of stars in some of the fields, this variation could also result from sampling effects or constraints in the observational design.
@@
Examining the star formation history of the SMC (Harris & Zaritsky, 2004, see their figures 3 and 6) we can see that the SMC has had a relatively constant star formation rate (SFR) from 2.5 Gyr until approximately 40 Myr ago, when the SFR started to decrease. All BLOeM fields have experience star-forming episodes in the last 40 to 10 Myr according to the SFR maps of Harris & Zaritsky (2004). Fields 6, 7, 8 are located in regions that experienced a strong SFR 10 Myr ago, but not in the last 6 to 4 Myr, whereas field 3 can be associated with strong star formation at 10 and 6 Myr, which is consistent with our suggestion that field 3 might contain a slightly younger population from the spectral types distribution analysis. Other fields that show a high SFR in the last 4 Myr from the maps of Harris & Zaritsky (2004) are 1 and 4, containing NGC 371 and NGC 346 respectively. We note however that our sample excludes O-type and supergiant stars, so while the presence (or absence) of early B stars can give a qualitative sense of recent star formation, it does not capture the youngest or most evolved populations. Thus, care must be taken when associating these stars alone with specific star-formation epochs. A quantitative analysis of the BLOeM stars by Bestenlehner et al. (in prep.) finds median ages for fields 1–8 in the range of 8–13 Myr, consistent with our identification of star-formation activity in the last 10 Myr, but reflecting a wider range of masses and evolutionary stages than those considered here.

### v4

@@
Binarity at LOw Metallicity (BLOeM):††thanks: Based on observations collected at the European Southern Observatory under ESO programme ID 112.25W2.
Early B-type stars with initial masses between 8 and 15 are frequently found in multiple systems, as is evidenced by multi-epoch spectroscopic campaigns in the Milky Way and the Large Magellanic Cloud (LMC). Previous studies have shown no strong metallicity dependence in the close-binary ( au) fraction or orbital-period distributions between the Milky Way’s solar metallicity () and that of the LMC (). However, similar analyses for a large sample of massive stars in more metal-poor environments are still scarce. We focus on 309 early B-type stars (luminosity classes III–V) from the Binarity at LOw Metallicity (BLOeM) campaign, which targeted nearly 1000 massive stars in the Small Magellanic Cloud (SMC, ) using VLT/FLAMES multi-epoch spectroscopy. By applying binary detection criteria consistent with previous works on Galactic and LMC samples, we identify 153 stars (91 SB1, 59 SB2, 3 SB3) exhibiting significant radial-velocity (RV) variations, resulting in an observed multiplicity fraction of % . Using Monte Carlo simulations to account for observational biases, we infer an intrinsic close-binary fraction of % . This fraction reduces to % when increasing our RV threshold from 20 to 80 km s-1; however, an independent Markov chain Monte Carlo analysis of the peak-to-peak RV distribution () confirms a high multiplicity fraction of % . These findings suggest a possible anti-correlation between metallicity and the fraction of close B-type binaries, with the SMC multiplicity fraction significantly exceeding previous measurements in the LMC and the Galaxy. The enhanced fraction of close binaries at SMC’s low metallicity may have broad implications for massive-star evolution in the early Universe. More frequent mass transfer and envelope stripping could boost the production of exotic transients, stripped supernovae, gravitational-wave progenitors, and sustained UV ionising flux, potentially affecting cosmic reionisation. Theoretical predictions of binary evolution under metal-poor conditions will provide a key test of our results.
@@
Massive stars go through complex evolutionary scenarios that are significantly influenced by their high multiplicity fraction (Mason et al., 2009; Chini et al., 2012; Sana et al., 2014; Kobulnicky et al., 2014; Dunstall et al., 2015; Moe & Di Stefano, 2017; Ramírez-Tannus et al., 2024) and frequent binary interactions (Sana et al., 2012; de Mink et al., 2014). Spectroscopic studies in the Galaxy and the Large Magellanic Cloud (LMC) consistently indicate that more than 50% of O- and B-type stars are part of binary or multiple systems within the orbital period range where interactions are expected ( d). Including longer orbital periods, the overall multiplicity fraction can exceed 90% for the O-type stars (Sota et al., 2014; Bordier et al., 2022). Sana et al. (2012) found an intrinsic multiplicity fraction (after correction for observational biases) of % for a sample of 71 O-type stars in young Galactic clusters. In the Cygnus OB2 association, Kobulnicky et al. (2014, and references therein) presented orbital solutions for 48 massive stars with spectral types ranging from O5 to B2, resulting in an intrinsic multiplicity fraction near 55% from a sample of 45 O-type and 83 B-type stars. More recently, Banyard et al. (2022) studied the B-type content of the young Galactic cluster NGC 6231, determining % from 80 stars covering the entire B-type spectral type range (B0-B9). While further multiplicity studies have been conducted in the Milky Way (MW), many have not been corrected for observational biases (e.g. Barbá et al., 2017; Berlanas et al., 2020; Ritchie et al., 2022), have targeted narrow spectral ranges (Abt et al., 1990), focused on peculiar types (e.g. chemically peculiar stars; Schöller et al., 2010), or addressed larger separations (Mason et al., 2009; Sana et al., 2014). These factors hinder direct comparison of multiplicity fractions due to the different observational strategies and biases.
@@
In the 30 Doradus (30 Dor) region of the LMC (), the VLT-FLAMES Tarantula Survey (VFTS; Evans et al., 2011) performed multiplicity studies on 360 O-type (Sana et al., 2013) and 408 B-type stars (Dunstall et al., 2015, of which 361 were main-sequence (MS) and giant stars), resulting in binary fractions111Throughout this paper, we use the terms multiplicity fraction and binary fraction interchangeably, and do not include higher-order multiples unless explicitly stated. of % and % , respectively. The consistency across these studies suggests that the multiplicity properties of massive stars may be independent of metallicity. Furthermore, the distribution of orbital periods is remarkably similar across most of these samples (Almeida et al., 2017; Villaseñor et al., 2021; Banyard et al., 2022), suggesting that O- and B-type stars in multiple systems are formed with similar efficiency and orbital configurations within the metallicity ranges studied so far, emphasising the need for multiplicity studies of massive stars in lower metallicity environments.
@@
In the case of solar-type binaries, a strong anti-correlation between binary fraction and metallicity has been found (Badenes et al., 2018; Moe et al., 2019; Offner et al., 2023). Moe et al. (2019) computed intrinsic binary fractions for five different observed samples comprising spectroscopic and eclipsing binaries (EBs), and found a combined close binary fraction that drops from % at [Fe/H] to % at [Fe/H]. This stands in stark contrast to the findings for OB-type stars (see also Moe & Di Stefano, 2013) and also to solar-type stars in wider orbits ( au; El-Badry & Rix, 2019). Both Moe et al. (2019) and El-Badry & Rix (2019) agree that the formation of wide binaries is governed by turbulent core fragmentation which is independent of metallicity, but, in the case of close binaries, the main formation mechanism is disc fragmentation up to separations of au. Moe et al. (2019) suggest that the difference between the sharp increase in the binary fraction of solar-type stars towards low metallicities and the close-to-constant fraction observed in massive stars lies in the disc of massive protostars. At solar metallicity, these discs are already highly unstable and have a high probability of fragmentation, leaving little room to further increase multiplicity at reduced metallicities.
@@
The findings in the MW and LMC suggest a universality in the multiplicity properties of massive stars, and motivate further investigation on whether trends persist at lower metallicity. Since the LMC metallicity is only half solar, it is essential to perform multiplicity studies of massive stars at even lower metallicity. In this context, the Small Magellanic Cloud (SMC) is a unique environment due to its proximity (62 kpc; Graczyk et al., 2020), low metallicity () representative of the high-redshift universe (Nakajima et al., 2023), and high content of massive stars (Evans et al., 2004). Furthermore, theoretical predictions indicate that the winds of massive stars at SMC metallicity are strongly reduced (Vink et al., 2001; Björklund et al., 2021), with empirical evidence showing that they might be even weaker than predicted (Ramachandran et al., 2019; Rickard et al., 2022). If there is a connection between metallicity and the multiplicity properties of massive stars, the SMC offers a prime environment for testing the universality of these properties. For other low-metallicity galaxies in the Local Group, resolving a significant sample of massive stars with the current generation of telescopes is not possible due to their greater distances.
@@
The Binarity at Low Metallicity (BLOeM; Shenar et al., 2024) campaign addresses this knowledge gap by focusing on the multiplicity properties of massive stars () within the SMC. BLOeM has observed 929 stars in eight different regions of the SMC. The sample has been divided into subsamples based on their spectral types, emission properties, and luminosity classes, enabling an interpretation of the results within the context of each subsample. The five samples are: the O-type stars (Sana et al., 2024); the B-type supergiants (spectral types B0-B3, Britavskiy et al., 2025); the B-type dwarfs and giants (this work); the cooler supergiants (spectral types later than B3, Patrick et al., 2025); and the emission-line stars (Oe/Be, Bodensteiner et al., 2025).
@@
In this study, we present the first detailed analysis of the multiplicity properties of a large sample of early B-type dwarf and giant stars at the low metallicity of the SMC. In Sect. 2, we introduce the sample of B-type stars and their main properties. Section 3 details our methodology to measure radial velocities (RVs) and our criteria to classify spectroscopic binaries. Our results on the multiplicity properties are presented in Sect. 4, and our search for periodicity is described in Sect. 5. We discuss our results in a broader context in Sect. 6, and finally provide our conclusions in Sect. 7.
@@ 2 The sample
The B-type dwarfs and giants sample comprises 309 stars with spectral types ranging from B0 to B2.5 and luminosity classes V to III222We note that O-type primaries with B-type companions are part of the O-type sample (Sana et al., 2024).. Spectral types for the full sample were presented in Shenar et al. (2024). Specifically, our sample consists of all early B-type “non-supergiant” objects, including MS (class V), subgiant (class IV), and relatively unevolved giant stars (class III). We expect our sample to primarily consists of MS stars undergoing core-hydrogen burning and stars that have just recently exhausted hydrogen in their cores. Bright giants and supergiants (classes II and I) form a separate early B-type supergiants subsample, which is the focus of Britavskiy et al. (2025). Hereafter, we refer to our sample as the III/V sample. This classification aligns with that used for the 30 Dor B-type stars, where Dunstall et al. (2015) defined “unevolved” stars as those with surface gravity dex and analysed their multiplicity properties separately from the supergiant sample. However, unlike the 30 Dor unevolved sample, our sample is mainly composed of giant stars (see Sect. 4.2). In the case of Be stars, we assume that most of them are binary interaction products (Bodensteiner et al., 2020b; Dallas et al., 2022; Dodd et al., 2024), therefore they will be the subject of a separate work (Bodensteiner et al., 2025).
BLOeM has currently obtained and reduced nine epochs from the first semester of observations using the FLAMES instrument (Fibre Large Array Multi Element Spectrograph; Pasquini et al., 2002) in GIRAFFE mode () on the UT2 unit telescope of the Very Large Telescope (VLT) at the European Southern Observatory (ESO) in Chile. For ten stars in our sample the number of epochs is below nine, fluctuating between three and seven epochs, and in five additional stars one of the epochs has too low a signal-to-noise ratio (), thus only eight epochs were used. Additional epochs are expected by the end of 2024, with a total of 25 epochs planned by the conclusion of the campaign in September 2025. For a complete overview of the sample selection, data reduction, and spectral classification, see Shenar et al. (2024).
@@
Each field’s central coordinates and the modified Julian dates (MJDs) per observation are provided by Shenar et al. (2024, Table A.1). Here, we present the total baseline of observations and minimum cadence for each field in Table 1. We have a mean observational baseline of 45 d and a minimum cadence of 1 to 2 d. This cadence allows us to constrain most short-period orbits, except for unevolved contact binaries with periods of less than d. However, these are likely to be EBs and therefore detectable by the Optical Gravitational Lensing Experiment (OGLE; Pawlak et al., 2016) through their light curves (see Sect. 5.1).
@@
The distribution of spectral types per field is shown in Fig. 1. The overall distribution is very similar to that found by Ramachandran et al. (2019) in the supergiant shell SMC-SGS1 in the Wing of the SMC, a region that was not covered by BLOeM. Our sample is dominated by spectral types B1.5 and B2, corresponding roughly to masses between 9 and 12. There is a good range of spectral types across different fields, but in Fields 7 and 8, the number of B1.5 and B2 seem to be overabundant with respect to the earliest stars (B0–B0.2), potentially suggesting an older population. We computed the ratio of latest-to-earliest type stars for these spectral types (i.e. (B1.5–B2)/(B0–B0.2)), finding ratios of 4.4 and 18.0 for Fields 7 and 8 respectively, and a mean of 2.9 for the other six fields. Performing a Z-test returns that these differences are significant, with a z-score of 5 (Field 7) and 49 (Field 8) and p-values of (Field 7) and 0.0 (Field 8). This is consistent with Fields 7 and 8 being areas without prominent star-forming regions (see Fig. 19 and details in the main text). In contrast, Field 3, which exhibits several H ii regions, has the lowest ratio of latest-to-earliest stars with 1.5. We note, however, that an uncertainty of one spectral subtype should be expected. Such uncertainty could at least partially explain the lack of B0.2 and B0.5 stars. These are stars that show weak He ii, and since He ii 4686 (stronger than He ii 4542) is not covered by the FLAMES LR02 setup, they are more difficult to classify. Given the small number of stars in some of the fields, this variation could also result from sampling effects or constraints in the observational design.
@@
Examining the star formation history of the SMC (Harris & Zaritsky, 2004, see their figures 3 and 6) we can see that the SMC has had a relatively constant star formation rate (SFR) from 2.5 Gyr until approximately 40 Myr ago, when the SFR started to decrease. All BLOeM fields have experienced star-forming episodes in the last 40 to 10 Myr according to the SFR maps of Harris & Zaritsky (2004). Fields 6, 7, 8 are located in regions that experienced a strong SFR 10 Myr ago, but not in the last 6 to 4 Myr, whereas Field 3 can be associated with strong star formation at 10 and 6 Myr, which is consistent with our suggestion that Field 3 might contain a slightly younger population from the spectral types distribution analysis. Other fields that show a high SFR in the last 4 Myr from the maps of Harris & Zaritsky (2004) are 1 and 4, containing NGC 371 and NGC 346 respectively. We note however that our sample excludes O-type and supergiant stars, so while the presence (or absence) of early B stars can give a qualitative sense of recent star formation, it does not capture the youngest or most evolved populations. Thus, care must be taken when associating these stars alone with specific star-formation epochs. A quantitative analysis of the BLOeM stars by Bestenlehner et al. (in prep.) finds median ages for Fields 1–8 in the range of 8–13 Myr, consistent with our identification of star-formation activity in the last 10 Myr, but reflecting a wider range of masses and evolutionary stages than those considered here.